A coil is joined in series with a pure resistor of resistance 80002 across a 100V, 50 Hz supply. The reading on a voltmeter across the coil is 45V and across the resistor is 80V. Determine the inductance and resistance of the coil. (ANS: 98.43812, 439.112 (1.398H)).


1
Expert's answer
2021-05-03T07:11:13-0400

Explanations & calculations


  • Consider the phasor diagram attached


  • The current flowing in the circuit can be calculated using the given information for the pure resistor.

"\qquad\qquad
\begin{aligned}
\small V&=\small iR\\
\small I_{total}&=\small \frac{80V}{80002\Omega}=\bold{9.99\times10^{-4}A}
\end{aligned}"


  • For this value of current, the given results are not received but for a value of "\small R=800\Omega" , the given results are obtained.

"\qquad\qquad
\begin{aligned}
\small I&=\small \frac{80V}{800\Omega}=0.1A
\end{aligned}"


  • To find the angle by which the coil voltage leads the circuit current, apply the parallelogram theorem.

"\qquad\qquad
\begin{aligned}
\small R&=\small \sqrt{P^2+Q^2+2PQ\cos\theta}\\
\small 100\Omega&=\small \sqrt{80^2+45^2+2(80)(45)\cos\theta}\\
\small \theta&=\small 77.36437^0
\end{aligned}"


  • Then considering the embedded sub-phasor diagram of the coil in the main diagram,


Voltage dropped for the coil resistance,

"\qquad\qquad
\begin{aligned}
\small V_{R,C}&=\small 45V\times\cos\theta=45\cos(77.36437)\\
&=\small 9.84375V
\end{aligned}"

Since the same current flows through this resistance,

"\qquad\qquad
\begin{aligned}
\small R_{coil}&=\small \frac{9.84375V}{0.1A}=\bold{98.4375\,\Omega}
\end{aligned}"


Voltage dropped for the inductance of the coil,

"\qquad\qquad
\begin{aligned}
\small V_{X,C}&=\small 45V\times\sin\theta=45\times\sin(77.36437)\\
&=\small 43.910V
\end{aligned}"

For a coil of inductance L,

"\qquad\qquad
\begin{aligned}
\small V&=\small 2\pi fL.i
\end{aligned}"

Using this, the inductance can be calculated (also the same current flows here)

"\qquad\qquad
\begin{aligned}
\small 43.910V&=\small 2\pi\times50Hz\times L\times0.1A\\
\small L&=\small \bold{1.3977H}
\end{aligned}"

Impedance of the coil if needed

"\qquad\qquad
\begin{aligned}
\small Z_{coil}&=\small 2\times\pi\times50Hz\times1.3977H=439.1004\Omega
\end{aligned}"